### v3

@@ 5.2.2 Human Evaluation
To validate the explanation results generated by the GPT-4 model, we conducted an online study with 17 participants with diverse demographics. The age distribution ranged from 18 to 55 years, with the majority (47.06%) in the 26-35 age group, followed by 35.29% in the 18-25 age group, 11.76% in the 36-45 age group, and 5.88% in the 46-55 age group. In terms of gender, 62.5% were male and 37.5% were female. Participants’ highest level of education varied, with 37.5% holding a PhD, 37.5% having a Master’s Degree, and 25% possessing a Bachelor’s Degree. Regarding expertise in explainability for autonomous agents, 37.5% were beginners, 31.25% had intermediate expertise, 12.5% were advanced, and 18.75% had no experience.
@@
The evaluation process involved three distinct questionnaires, one corresponding to each scenario. Participants assessed 16 explanations per scenario, resulting in a total of 48 explanations evaluated by each individual. Each explanation was rated using five questions derived from the metrics outlined in Table 4, based on the Explanation Satisfaction Scale proposed by Hoffman et al. in [102]. This scale represents a widely recognized framework for measuring human responses to XAI systems. To ensure the evaluation aligned with human preferences, we used a 5-point Likert scale. Answers in this approach are distributed in 1 (I disagree strongly), 2 (I disagree somewhat), 3 (I’m neutral about it), 4 (I agree somewhat), and 5 (I agree strongly).
@@
In addition to the explanation evaluations and optional demographic information, the questionnaire provided general instructions describing the robot’s task and included a map for each scenario. The questionnaires were distributed via a mailing list, and participants were informed about the anonymity and research purposes of the study. Consent was obtained upon submission of their answers. Participation was voluntary, which made the process cost-effective but also introduced potential drawbacks, such as longer response times and biases that may arise from factors like age, education level, or familiarity with autonomous system explainability approaches.
@@
Statistical analysis is essential in human evaluations to ensure rigor and reliability, particularly when interpreting subjective judgments and addressing variability among participants. Likert-scale data, which is ordinal, requires careful statistical treatment to identify significant trends and differences while avoiding misinterpretation of the results. By applying appropriate statistical methods, we strengthen the credibility of our conclusions and provide a robust evaluation of the LLM’s results across scenarios.
@@
To ensure the reliability and robustness of the evaluation, a series of statistical tests were applied. Reliability analysis of the scale criteria was conducted using Cronbach’s Alpha, a robust metric for assessing internal consistency and ensuring that the scale reliably measures the intended constructs, with a commonly accepted threshold for reliability being $\alpha>0.7$ [103, 104].
@@
To further explore the relationships among variables in the dataset, correlation analysis was performed using Pearson correlation matrices, which provide insights into the strength and direction of linear relationships between variables [105].
@@
For inferential analysis, a normality check was conducted using the Shapiro-Wilk test to determine whether the data met the assumptions required for parametric tests [106]. Since the data violated these assumptions, the non-parametric Friedman test was employed [107]. The Friedman test is particularly suited for analyzing repeated measures on the same participants, as it does not require data to follow a normal distribution [108]. This test assessed whether significant differences existed in participant responses across the three scenarios, offering a reliable framework for understanding the impact of different conditions on the perceived quality and effectiveness of explanations.
@@
Finally, post-hoc analysis using the Nemenyi test was performed to examine participant behavior in specific explanations where significant differences were observed, offering a deeper understanding of the underlying patterns in their responses. The statistical analyses were developed through Python scripts (version 3.10.12), utilizing libraries such as pingouin (version 0.5.5) for Cronbach’s Alpha, scipy.stats (version 1.14.1) for the Shapiro-Wilk and Friedman tests, and pandas (version 2.3.3) for data manipulation. This approach ensured accuracy and reproducibility in the execution of the tests.
Table 4: Questions and metrics used for human evaluation.
@@
Usefulness The explanation is useful to my goals
5.2.3 Categories and Evaluation Questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into the different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the BT. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [109].
The set of categories, their descriptions and the questions included in each one are described next.
@@ 5.2.3 Categories and Evaluation Questions
These questions provide an overview of the navigation process captured in the Rosbag file. Understanding what has happened and how the navigation task has proceeded provides a high-level context that sets the stage for more detailed questions in later categories. This category comprises the following questions:
@@
These questions delve into the specifics of trajectory planning, a critical aspect of navigation. They address whether the robot has re-planned its trajectory, the reasons behind this behavior, and whether it has encountered any obstacles. These details are necessary for understanding how the robot dynamically adjusts its path. The questions included in this category are:
@@
These questions examine the specifics of goals and locations. Understanding the location of goals, the robot’s linear velocity during navigation, its initial position and orientation, and the velocity after receiving a goal contribute to a detailed understanding of the robot’s movements and performance. Questions included in this category are:
@@
    What is the initial position and orientation of the robot?
@@
    Which Behavior Tree’s nodes were used during navigation?
6 Experimental Results and Discussion
This section summarizes the experiments’ results involving both the black box-like engine and the explainability component when using our solution in the previously depicted assessment scenarios.
6.1 Message Loss Rates
Figure 7 shows the percentage of lost messages in the Rosbag file for experiments detailed in Section 5, capturing the topics outlined in Table 2, attending to their publication rates and to the approaches selected to store integrity proofs in blockchain. The results across the three scenarios demonstrate a consistent trend, showing uniform message loss patterns in the creation of immutable Rosbag files.
Choosing to store a single message hash for every 100 messages results in a loss of less than 5% compared to avoiding anti-tampering measures in the Rosbag recording process, even for topics with rates exceeding 300 Hz. This approach stands out as the most efficient option for integrating anti-tampering techniques into our accountability solution.
Alternative strategies, such as saving one proof every 50 messages or adjusting proof intervals based on topic rates, achieve losses below 10% for each topic. This is particularly noteworthy, ensuring the integrity of messages based on their significance in a given scenario or task.
Introducing an integrity proof every 25 messages leads to an overall reduction in message recording, especially evident when topic rates surpass 100 Hz. Storing one hash every 10 messages could lead to losses exceeding 30% for topics with rates higher or around 100 Hz. Beyond the consistent linear trend of message loss with high topic rates, a slight deviation is observed in this approach when topics fall below 50 Hz. This variation signifies the influence on message recording when computing hash functions for messages nearing 1 MB in size, as observed in messages from the /camera/image_raw topic, given that the computational complexity of common hash functions, such as SHA-256, increases with input size.
@@ 6.1 Message Loss Rates
6.2 System Benchmarking
The system load or, the running thread demand on the system as the average number of running and idle threads, is shown in Figure 8. The average of 1 minute is higher than the average of 5 or 15 minutes in all scenarios, showing a load increase. However, the values of this metric do not exceed the CPU count, which is 6 in our experimental setup, indicating no performance issues or resource saturation. The inclusion of runnable and uninterruptible tasks in this metric means that the load average may increase due to a disk I/O workload, not just CPU demand, giving a more general view of system overhead. Other more specific metrics are described next to clarify the possible existence of a bottleneck in our approach.
@@ 6.2 System Benchmarking
Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity. Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach does not have a remarkable impact on the system performance.
Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact. In addition, all the approaches that perform integrity-proof storage on the blockchain achieve lower or similar results to the approach avoiding anti-tampering operations.
@@
6.3 Correctness Evaluation Using LLM-as-a-judge
Figure 10 shows the results of the correctness of RAG evaluation, presenting both the overall results and assessments based on predefined categories. The results demonstrate consistency across the three scenarios, getting a significant number of correct answers, even though there has not been a prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@ 6.3 Correctness Evaluation Using LLM-as-a-judge
Regarding individual categories, the correctness in the category ‘Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question "What has happened in this ROS 2 log regarding navigation?" initially yielded a correctness below 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failures of Nav2 Behavior Tree nodes, not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized BTs information, other times focused on locations or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension "Describe the sequence of events regarding navigation goals and include information about each of them." in Question 1 significantly increased correctness in the first category, achieving results around and exceeding 75% in all three scenarios. This fact underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
@@
The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning. The model can determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or other circumstance. Scenario 1 and Scenario 2 also exhibit relatively high correctness in questions related to goal completion and task status. However, the model encounters challenges in Scenario 3, occasionally failing to identify whether a navigation that starts correctly concludes with an abortion or cancellation. This issue may be attributed to misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
@@
Correctness in identifying specific targets and locations lacks some consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third. In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, as the model misaligns the queried position with the one assumed by the robot when starting the navigation process. This fact may also be attributed to deficiencies in the retrieval process. In Scenario 1, understanding the sequence of events or the order of goals is a challenge for the model, leading to inaccuracies in specifying the second location or goal. Nevertheless, this challenge is not present in Scenario 3 and is absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
@@
Through questions included in the ‘Nav2 Behavior Tree and Node Status’ category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [110]. While the model excels in identifying BT nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the criteria correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 BT nodes not present in the other scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 BT node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the complexities of the recovery actions and the dynamic nature of BT executions during navigation recovery introduce complexities that the model may not fully generalize from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
6.4 Criteria-based Evaluation Using LLM-as-a-judge
@@
The language model’s adaptability to challenges like obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
@@
Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness improve, Scenario 1, which involves direct navigation to three goals, scores higher in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
In the category ‘Trajectory Planning and Re-planning’, questions are effectively addressed across all three scenarios, showing scores above 8 in every metric, even when obstacles or goal cancellations are introduced. These results highlight the system’s capability to manage trajectory-related queries effectively despite challenges.
Regarding the category ‘Goal Completion and Navigation Task Status’, Scenarios 1 and 2 effectively convey information about goal completion and navigation task progress. Nevertheless, slightly lower scores in Scenario 3 indicate challenges in effectively communicating completion status when a goal is aborted.
@@
The results regarding questions included in ‘Nav2 Behavior Tree and Node Status’ category demonstrate effectiveness in communication across all scenarios and metrics. This emphasizes the potential for BTs to complement language models in explainability, offering insight into the decision-making processes during navigation tasks.
In summary, although results can be improved in accuracy for general questions, the language model shows adaptability across different navigation scenarios, including those with obstacles and task changes. The consistency in conveying information related to Nav2 BTs suggests a solid understanding of underlying navigation processes. The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario. Nevertheless, while evaluation using LLMs excels in scalability and efficient resource usage, it is still recommended to test a subset of questions with human evaluators. This ensures a deeper understanding of how well the approach aligns with human expectations and effectively communicates, particularly in scenarios involving complex or nuanced details.
6.5 Human Evaluation of LLM-generated Explanations
This section presents an analysis of the results from the human evaluation conducted with 17 participants, as stated in Section 5.2.2.
6.5.1 Distribution Analysis
Figure 13 depicts the distribution of Likert scale responses for the five evaluation criteria across the three scenarios, showing the frequency of each response value. The highest ratings predominate across all scenarios and criteria, suggesting that participants generally found the explanations satisfactory. However, Scenario 2 and Scenario 3 exhibit greater variability, with an increase in mid-range scores for certain criteria, such as Usefulness. These trends are particularly noticeable in broad questions, such as those in Category 1 regarding the navigation process overview, or when the model’s answers do not align with the ground truth. In such cases, participant opinions diverge significantly, even in the presence of factual inaccuracies, indicating a potential bias in how the utility is perceived when discrepancies with the ground truth arise.

### v4

@@ V-B2 Human Evaluation
To validate the explanation results generated by the GPT-4 model, we conducted an online study with 17 participants with diverse demographics. The age distribution ranged from 18 to 55 years, with the majority (47.06%) in the 26-35 age group, followed by 35.29% in the 18-25 age group, 11.76% in the 36-45 age group, and 5.88% in the 46-55 age group. In terms of gender, 62.5% were male and 37.5% were female. The participants’ highest level of education varied, with 37.5% holding a PhD, 37.5% having a master’s degree, and 25% possessing a bachelor’s degree. Regarding expertise in explainability of autonomous agents, 37.5% were beginners, 31.25% had intermediate expertise, 12.5% were advanced, and 18.75% had no experience.
@@
The evaluation process involved three questionnaires, with one corresponding to each scenario. The participants assessed 16 explanations per scenario, resulting in a total of 48 explanations evaluated by each individual. Each explanation was rated using five questions derived from the metrics outlined in Table IV, based on the Explanation Satisfaction Scale proposed by Hoffman et al. [95]. This scale is a widely recognized framework for measuring human responses to XAI systems. To ensure that the evaluation aligned with human preferences, we used a 5-point Likert scale. Answers in this approach are distributed in 1 (I disagree strongly), 2 (I disagree somewhat), 3 (I’m neutral about it), 4 (I agree somewhat), and 5 (I agree strongly).
@@
In addition to the explanation evaluations and optional demographic information, the questionnaire provided general instructions describing the robot’s task and included a map for each scenario. The questionnaires were distributed via a mail list, and participants were informed about the anonymity and research purposes of the study. Consent was obtained upon submission of their answers. Participation was voluntary, which made the process cost-effective but also introduced potential drawbacks, such as longer response times and biases that may arise from factors such as age, education level, or familiarity with autonomous system explainability approaches.
@@
Statistical analysis is essential in human evaluations to ensure rigor and reliability, particularly when interpreting subjective judgments and addressing variability among participants. Likert-scale data, which is ordinal, requires careful statistical treatment to identify significant trends and differences while avoiding misinterpretation of the results. By applying appropriate statistical methods, we strengthen the credibility of our conclusions and provide a robust evaluation of LLM’s results across scenarios.
@@
A series of statistical tests was applied to ensure the reliability and robustness of the evaluation. Reliability analysis of the scale criteria was conducted using Cronbach’s Alpha, a robust metric for assessing internal consistency and ensuring that the scale reliably measures the intended constructs, with a commonly accepted threshold for reliability of $\alpha>0.7$ [96].
@@
To further explore the relationships among variables in the dataset, correlation analysis was performed using Pearson correlation matrices, which provide insights into the strength and direction of the linear relationships between variables [97].
@@
For inferential analysis, a normality check was conducted using the Shapiro-Wilk test to determine whether the data met the assumptions required for parametric tests. As the data violated these assumptions, the non-parametric Friedman test was employed [98]. The Friedman test is particularly suited for analyzing repeated measures on the same participants, as it does not require data to follow a normal distribution. This test assessed whether significant differences existed in participant responses across the three scenarios, offering a reliable framework for understanding the impact of different conditions on the perceived quality and effectiveness of explanations.
@@
Finally, a post-hoc analysis using the Nemenyi test was performed to examine participant behavior in specific explanations where significant differences were observed, offering a deeper understanding of the underlying patterns in their responses. Statistical analyses were performed through Python scripts (version 3.10.12), utilizing libraries such as pingouin (version 0.5.5) for Cronbach’s Alpha, scipy.stats (version 1.14.1) for the Shapiro-Wilk and Friedman tests, and pandas (version 2.3.3) for data manipulation. This approach ensured accuracy and reproducibility during the text execution.
TABLE IV: Questions and metrics used for human evaluation.
@@
Usefulness The explanation is useful to my goals.
V-B3 Categories and Evaluation Questions
To evaluate our explainability proposal in the context of ROS 2 message interpretation, we defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into the different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we organized them into five categories. These categories range from high-level overviews of navigation to detailed inquiries regarding trajectory planning, goal completion, and BT functioning. These categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [99].
The set of categories, their descriptions and questions included in each one are described next.
@@ V-B3 Categories and Evaluation Questions
These questions provide an overview of the navigation process captured at the Rosbag file. Understanding what has occurred and how the navigation task has proceeded provides a high-level context that sets the stage for more detailed questions in later categories. This category comprises the following questions:
@@
These questions delve into the specifics of trajectory planning, which is a critical aspect of navigation. They address whether the robot has re-planned its trajectory, the reasons behind this behavior, and whether it has encountered any obstacles. These details are necessary to understand how a robot dynamically adjusts its path. The questions included in this category are:
@@
These questions examine the specifics of the goals and locations. Understanding the location of goals, the robot’s linear velocity during navigation, its initial position and orientation, and the velocity after receiving a goal contribute to a detailed understanding of the robot’s movements and performance. The questions included in this category are:
@@
    What are the initial position and orientation of the robot?
@@
    Which Behavior Tree nodes were used during navigation?
VI Experimental Results and Discussion
This section summarizes the results of the experiments involving both the black box-like engine and explainability component when using our solution in the previously depicted assessment scenarios.
VI-A Message Loss Rates
Figure 7 shows the percentage of lost messages in the Rosbag file for the experiments detailed in Section V, capturing the topics outlined in Table II, attending to their publication rates, and to the approaches selected to store integrity proofs in the blockchain. The results across the three scenarios demonstrate a consistent trend, showing uniform message loss patterns in the creation of immutable Rosbag files.
Choosing to store a single message hash for every 100 messages results in a loss of less than 5% compared with avoiding anti-tampering measures in the Rosbag recording process, even for topics with rates exceeding 300 Hz. This approach is the most efficient option for integrating anti-tampering techniques into our accountability solution.
Alternative strategies, such as saving one proof every 50 messages or adjusting proof intervals based on topic rates, achieve losses of less than 10% for each topic. This is particularly noteworthy, as it ensures the integrity of messages based on their significance in a given scenario or task.
Introducing an integrity proof every 25 messages leads to an overall reduction in message recording, particularly when topic rates surpass 100 Hz. Storing one hash every 10 messages could lead to losses exceeding 30% for topics with rates higher than or around 100 Hz. Beyond the consistent linear trend of message loss with high topic rates, a slight deviation is observed in this approach when topics fall below 50 Hz. This variation signifies the influence on message recording when computing hash functions for messages nearing 1 MB in size, as observed in messages from the /camera/image_raw topic, given that the computational complexity of common hash functions, such as SHA-256, increases with input size.
@@ VI-A Message Loss Rates
VI-B System Benchmarking
The system load or, running thread demand on the system as the average number of running and idle threads, is shown in Figure 8. The average of 1 minute is higher than the average of 5 or 15 minutes in all scenarios, showing a load increase. However, the values of this metric do not exceed the CPU count, which is 6 in our experimental setup, indicating no performance issues or resource saturation. The inclusion of runnable and uninterruptible tasks in this metric means that the load average may increase due to a disk I/O workload, not just CPU demand, giving a more general view of the system overhead. Other more specific metrics are described next to clarify the possible existence of a bottleneck in the proposed approach.
@@ VI-B System Benchmarking
Regarding network traffic, approaches with higher process storage in the blockchain increase this metric, but do not exceed 0,20 MB/s, which is a very affordable quantity. Furthermore, approaches such as storing one hash every 100 messages nearly match the avoidance of storing any proof into the blockchain, indicating that including this approach does not have a remarkable impact on system performance.
Disk write operations do not form a bottleneck in system performance, with all approaches differing by no more than 0,40 MB/s, which is not a very significant impact. In addition, all approaches that perform integrity-proof storage on the blockchain achieve lower or similar results to the approach that avoids anti-tampering operations.
@@
VI-C Correctness Evaluation Using LLM-as-a-judge
Figure 10 shows the results of the correctness of the RAG evaluation, presenting both the overall results and the assessments based on predefined categories. The results demonstrate consistency across the three scenarios, obtaining a significant number of correct answers, even though there has not been prior tuning in the ROS 2 messages for autonomous agents’ navigation tasks. The slight percentage of incorrect answers remains within an acceptable threshold, given the specific context employed and the inclusion of general questions.
@@ VI-C Correctness Evaluation Using LLM-as-a-judge
Regarding individual categories, the correctness in the category ‘Navigation Process Overview’ is relatively consistent across all scenarios, indicating a good understanding of events and progress related to navigation. However, the broad nature of the question ”What has happened in this ROS 2 log regarding navigation?” initially yielded a correctness rate of less than 65% in the third scenario. In this instance, the cancellation of a goal resulted in the execution and failure of Nav2 Behavior Tree nodes, which were not included in the other two scenarios and focused on navigation recovery. Therefore, the incorporation in a dataset of information related to the cancellation of a goal results in LLM answers that usually diverge from the ground truth provided, as the model occasionally prioritized BTs information, other times focused on locations, or timestamps. This variability made it difficult to align with the ground truth provided, even if the obtained answers could be considered partially accurate. Introducing the prompt extension ”Describe the sequence of events regarding navigation goals and include information about each of them.” in Question 1 significantly increased the correctness in the first category, achieving results around and exceeding 75% in all three scenarios. This underscores the effectiveness of prompt engineering when tackling ambiguous or broad queries.
@@
The results also show consistently high correctness across all scenarios concerning questions related to trajectory planning and re-planning. The model can determine whether the modification of a predicted trajectory has been caused by the occurrence of an obstacle or by other circumstance. Scenarios 1 and 2 also exhibit relatively high correctness in questions related to goal completion and task status. However, the model encounters challenges in Scenario 3, occasionally failing to identify whether a navigation that starts correctly concludes with an abortion or cancellation. This issue may be attributed to a misalignment during the retrieval process, specifically in the selection of the best chunk associated with a particular goal associated with numerous embeddings within a single collection from the vector database.
@@
Correctness in identifying specific targets and locations lacks consistency, with results exceeding 50% in the first scenario, around 70% in the second and over 75% in the third. In particular, three scenarios encounter difficulties in identifying the initial position of the robot before starting the navigation, as the model misaligns the queried position with that assumed by the robot when starting the navigation process. This may also be attributed to deficiencies in the retrieval process. In Scenario 1, understanding the sequence of events or order of goals is a challenge for the model, leading to inaccuracies in specifying the second location or goal. Nevertheless, this challenge is not present in Scenario 3 and is absent or reduced in Scenario 2, demonstrating how differences in data distribution can introduce varied responses even when using specific and unambiguous questions.
@@
Through questions included in the ‘Nav2 Behavior Tree and Node Status’ category, we assessed the effectiveness of the model’s formal linguistic competence compared to its functional linguistic competence, a feature of present LLMs [100]. While the model excels in identifying BT nodes given its description, it faces challenges in pinpointing these components under specific conditions. As a result, the criteria correctness suffers a drop in Scenario 3, although overall results remain above 75%. In this scenario, the navigation recovery actions caused by the cancellation of the first goal introduced a challenge in the interpretation of the information by the LLM due to the execution of several Nav2 BT nodes not present in the other scenarios. Particularly, the mistakes stem from the model’s confusion in determining whether a previously failed Nav2 BT node has recovered or finished its execution, as well as identifying all nodes that were executed during navigation. This can be attributed to the fact that the nuanced behavior of recovery actions and the dynamic execution of BTs during navigation introduce difficulties that the model may not adequately capture from the training data, underscoring the need for additional fine-tuning in such scenario-specific contexts in case the obtained correctness needs to be improved.
VI-D Criteria-based Evaluation Using LLM-as-a-judge
@@
The adaptability of the language model to challenges such as obstacles and goal cancellations regarding the described categories, is shown in Figure 12.
@@
Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness improve, Scenario 1, which involves direct navigation to the three goals, scores higher in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
In the category ‘Trajectory Planning and Re-planning’, questions are effectively addressed across all three scenarios, showing scores above 8 in every metric, even when obstacles or goal cancellations are introduced. These results highlight the system’s capability to manage trajectory-related queries effectively despite these challenges.
Regarding the category ‘Goal Completion and Navigation Task Status’ category, Scenarios 1 and 2 effectively convey information about goal completion and navigation task progress. Nevertheless, the slightly lower scores in Scenario 3 indicate challenges in effectively communicating the completion status when a goal is aborted.
@@
The results regarding questions included in ‘Nav2 Behavior Tree and Node Status’ category demonstrate effectiveness in communication across all scenarios and metrics. This emphasizes the potential of BTs to complement language models in explainability, offering insight into decision-making processes during navigation tasks.
In summary, although the results can be improved in accuracy for general questions, the language model shows adaptability across different navigation scenarios, including those with obstacles and task changes. Consistency in conveying information related to Nav2 BTs suggests a solid understanding of underlying navigation processes. The model also performs well when providing information about trajectory planning, goal completion, and specifics about goals and locations, regardless of the specific challenges introduced in each scenario. Nevertheless, although evaluation using LLMs excels in scalability and efficient resource usage, it is still recommended to test a subset of questions with human evaluators. This ensures a deeper understanding of how well the approach aligns with human expectations and communicates effectively, particularly in scenarios involving complex or nuanced details.
VI-E Human Evaluation of LLM-generated Explanations
This section presents an analysis of the results of the human evaluation conducted with 17 participants, as stated in Section V-B2.
VI-E1 Distribution Analysis
Figure 13 depicts the distribution of Likert scale responses for the five evaluation criteria across the three scenarios, showing the frequency of each response value. The highest ratings predominate across all scenarios and criteria, suggesting that participants generally found the explanations to be satisfactory. However, Scenarios 2 and 3 exhibit greater variability, with an increase in mid-range scores for certain criteria, such as Usefulness. These trends are particularly noticeable in broad questions, such as those in Category 1 regarding the navigation process overview, or when the model’s answers do not align with the ground truth. In such cases, participant opinions diverge significantly, even in the presence of factual inaccuracies, indicating a potential bias in how utility is perceived when discrepancies with the ground truth arise.